OpenStudy (lena772):

A company produces accessories for smart phones and tablets. The profit on each smart phone case is $2 and the profit on each tablet case is $3. The company made a profit of $1,200 on the cases last month. The equation 2x + 3y = 1,200 represents the company's profit from cases last month, where x is the number of smart phone cases sold and y is the number of tablet cases sold. Change the equation into slope-intercept form. Identify the slope and y-intercept of the equation. Be sure to show all of your work.

OpenStudy (anonymous):

1200=2x+3y 0=2x+3y-1200 -3y=2x-1200 divide everything by (-3) y=(-2/3)x+400 m (slope)= -2/3 b (y intercept)= 400 :)

OpenStudy (lena772):

That's it?

OpenStudy (anonymous):

yep! to put an equation into slope intercept form, all you need to do is isolate y and you have your equation!

OpenStudy (lena772):

Thanks for the help